# On the concentration distribution in turbulent thermals

### v1

@@ 6 Discussion and conclusion
[Acknowledgements]LH thanks Daniel Lecoanet for helping implement Dedalus into the HPC of Université Grenoble Alpes and for the data provided for the large-resolution simulations. All the 3D simulations presented in this paper were performed using the GRICAD infrastructure (https://gricad.univ-grenoble-alpes.fr), supported by Grenoble research communities.
[…]
[Funding] This work was supported by the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant number 716429). ISTerre is part of Labex OSUG@2020 (ANR10 LABX56).
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
[Data availability statement]
[…]
[Author ORCIDs]L. Huguet, https://orcid.org/$0$-0002-5532-6302; V. Lherm, https://orcid.org/$0$-0001-5814-0637; R. Deguen, https://orcid.org/$0$-0002-6883-0655; T. Le Borgne, https://orcid.org/$0$-0001-9266-9139; J. Heyman, https://orcid.org/$0$-0002-0327-7924
@@ Appendix A Intensity-Concentration Calibrations
[…]
Figure 11 shows one of the calibrations of the light intensity $I$ as a function of the concentration $c$ of Rhodamine 6G. When $c\to 0$, the light intensity should increase linearly with the dye concentration and attenuate at high concentrations. This is indeed the case when $c\gtrsim 1~{}$\mathrm{m}\mathrm{g}\mathrm{/}\mathrm{L}$$. This attenuation could also be related to interactions with NaCl, leading to the formation of aggregates that reduce the fluorescence intensity. To account for this effect at high concentration, the intensity-concentration calibration thus employs a power law of the form $I=ac^{b}$ rather than a linear law.
Appendix B Supplementary figures at different Reynolds numbers
Figure 12 shows the evolution of the thermal for experiments and simulations for all Reynolds numbers. Standard deviation of the concentration field for the entire time series and each of the three experiments or two orthogonal plans of 3D simulations. Red dashed lines correspond to the evolution of the radius of the thermal with depth $r=\alpha z$, using $\alpha=0.18$.
Figure 12: Standard deviation of the concentration field for the entire time series and each of the three experiments or two orthogonal plans of 3D simulations. Red dashed lines correspond to the evolution of the radius of the thermal with depth $r=\alpha z$, using $\alpha=0.18$.
The appendix B also contains figures of the PDF of concentration for the different experiments and simulations presented in this manuscript. Figures 13 and 14, and 15 show the PDF of concentration for 3 Reynolds numbers. Figures 16, 17, and 18 are the PDFs of the concentration normalized by the mean of $c$. Figure 19,20, 21, and 22 are the cumulative PDFs of the $c/\langle c\rangle$ (see 8). Figure 23 shows the PDFs of the concentration normalized by the mean of $c$ for two simulations with the same Reynolds but with two different resolutions $256\times 256\times 512$ and $512\times 512\times 1024$.
Figure 13: PDF of $c$ as a function of $c$ and time $t/t_{g}$ with a given $Re=2012$ for one of the three experiments (a,c) ($Pe=5.04\times 10^{6}$) and a simulation (b,d) ($Pe=2.8\times 10^{3}$). The top and bottom rows are log-log and semi-log plots, respectively.
Figure 14: Same as figure 13 for $Re=4175$ ($Pe=1.04\times 10^{7}$ for experiments and $Pe=4.2\times 10^{3}$ for simulation ).
Figure 15: Same as figure 13 for $Re=4646$ ($Pe=1.16\times 10^{7}$ for experiments and $Pe=4.6\times 10^{3}$ for simulation ).
Figure 16: PDF of $c/(\langle c\rangle)$ as a function of $c/(\langle c\rangle)$ and time $t/t_{g}$ with a given $Re=2012$ for an experiment (a, c) ($Pe=5.04\times 10^{6}$) and a simulation (b, d) ($Pe=2.0\times 10^{3}$). The top and bottom rows are log-log and semi-log plots, respectively. The gray denotes the equation of the exponential distribution 24.
Figure 17: Same as figure 16 for $Re=4175$ ($Pe=1.04\times 10^{7}$ for experiments and $Pe=4.2\times 10^{3}$ for simulation ).
Figure 18: Same as figure 16 for $Re=4646$ ($Pe=1.16\times 10^{7}$ for experiments and $Pe=4.6\times 10^{3}$ for simulation ).
Figure 19: Cumulative PDF of $c/\langle c\rangle$ as a function of $c/\langle c\rangle$ and time $t/t_{g}$ for an experiment with $Re=3452$ (a, c) ($Pe=8.64\times 10^{6}$) and a simulation at the same $Re$ (b, d) ($Pe=3.4\times 10^{3}$). The top and bottom rows are log-log and linear-linear plots, respectively.
Figure 20: Same as figure 19 for $Re=2012$ ($Pe=5.04\times 10^{6}$ for experiments and $Pe=2.0\times 10^{3}$ for simulation ).
Figure 21: Same as figure 19 for $Re=4175$ ($Pe=1.04\times 10^{7}$ for experiments and $Pe=4.2\times 10^{3}$ for simulation ).
Figure 22: Same as figure 16 for $Re=4646$ ($Pe=1.16\times 10^{7}$ for experiments and $Pe=4.6\times 10^{3}$ for simulation ).
Figure 23: PDF of $c/(\langle c\rangle)$ as a function of $c/(\langle c\rangle)$ and time $t/t_{g}$ with a given $Re=4631$ for a simulation with a resolution of $256\times 256\times 512$ (a, c) and with a resolution of $512\times 512\times 1024$ (b, d). The top and bottom rows are log-log and semi-log plots, respectively. The gray denotes the equation of the exponential distribution 24.

### v2

@@ Acknowledgements.
LH thanks Daniel Lecoanet for helping implement Dedalus into the HPC of Université Grenoble Alpes and for the data provided for the large-resolution simulations. All the 3D simulations presented in this paper were performed using the GRICAD infrastructure (https://gricad.univ-grenoble-alpes.fr), supported by Grenoble research communities. This work was supported by the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (grant number 716429). ISTerre is part of Labex OSUG@2020 (ANR10 LABX56).
@@ Appendix A Intensity-Concentration Calibrations
Figure 11: Fluorescence light intensity as a function of Rhodamine 6G concentration in aqueous solutions of NaCl. The solid line corresponds to a power law of the form $I=ac^{b}$ fitted to the experimental data, with $a=(1.60\pm 0.01)\times${10}^{4}$\$ and $b=0.93\pm 0.01$. The dashed line corresponds to a power law of the form $I=ac$ fitted to the experimental data for $c\leq 1~$\mathrm{m}\mathrm{g}\mathrm{/}\mathrm{L}$\$, with $a=(1.63\pm 0.02)\times${10}^{4}$\$.
Fig. 11 shows one of the calibrations of the light intensity $I$ as a function of the concentration $c$ of Rhodamine 6G. When $c\to 0$, the light intensity should increase linearly with the dye concentration and attenuate at high concentrations. This is indeed the case when $c\gtrsim 1~$\mathrm{m}\mathrm{g}\mathrm{/}\mathrm{L}$\$. This attenuation could also be related to interactions with NaCl, leading to the formation of aggregates that reduce the fluorescence intensity. To account for this effect at high concentration, the intensity-concentration calibration thus employs a power law of the form $I=ac^{b}$ rather than a linear law.